OpenStudy (anonymous):

Write the linear equation below in slope-intercept form. Show the steps you used to arrive at your solution. 5x - 3y = 24

OpenStudy (anonymous):

Again, slope intercept form is y=mx+b so you have 5x - 3y = 24 first thing youre going to do is move the 5x to the other side do that by subtracting 5x -3y = 24 -5x -5x -3y = -5x + 24 Now we almost have it in s-i form, but we have to get rid of the 3 in front of the y, do this by diving it into every variable on the opposite side. -3y = -5x + 24 /-3 /-3 your answer is y = 5/3 -8

OpenStudy (anonymous):

thank yuh